OpenStudy (anonymous):

Find the 10th term of an arithmetic sequence if t1 = 2.1 and t4 = 1.83.

OpenStudy (amistre64):

2.1 + 3n = 1.83; solve for n

OpenStudy (amistre64):

gonna nail it down soon lol

OpenStudy (amistre64):

lets start over....

OpenStudy (amistre64):

2.1 +x +x +x = 1.83 3x = -.27 x = -.27/3 = -.09; check t1 = 2.1 t2 = 2.1 -.09 = 2.01 t3 = 2.01 - .09 = 1.92 t4 = 1.92 - .09 = 1.83 checks out

OpenStudy (amistre64):

t1 = t0 - .09 = 2.1 t0 = 2.1 + .09 = 2.19 tn = 2.19 - .09n; chk t4 = 2.19 -.09(4) = 1.83 ..good t10 = 2.19 - .09(10) = 1.29